OpenStudy (anonymous):

Use the distributive property to rewrite the expression (a+8)4

OpenStudy (raden):

(a+8)4 = 4 * a + 4 * 8 = ....

OpenStudy (anonymous):

ok

OpenStudy (raden):

so, what's ur answer ?

OpenStudy (anonymous):

4a+32

OpenStudy (raden):

smart ;)

OpenStudy (anonymous):

:)